OpenStudy (anonymous):

A system of equations is shown below: 5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

@nikato

OpenStudy (triciaal):

the first eqn is the same as x - y = 2 if you divide everything by 5

OpenStudy (anonymous):

5x-5y+5y=10+5y 5x=5y+10 ___________ 5 x=y+2

OpenStudy (triciaal):

x - y = 2 eqn 1 3x -2y = 2 eqn 2 multiply eqn 1 by -2 -2x +2y = -4 eqn 3 add eqn 2 and eqn 3 x = -2 put this value of x in either eqn to find y using eqn 1 -2 -y = 2 rearrange to get y = -4 x = -2 and y = -4 substitute in original eqn left hand side 5x -5y = 5(-2) - 5(-4) = -10 + 20 = 10 right hand side = 10 10 = 10

OpenStudy (triciaal):

@RosieF any questions?